# Distributional Reduction: Unifying Dimensionality Reduction and Clustering with Gromov-Wasserstein

### v2

@@ Appendix E Appendix of experimental section
  • E.1: implementation details, validation of hyperparameters, datasets and metrics.
  • E.2: comparison with COOT clustering.
  • E.3: complete scores on all datasets.
  • E.4: study homogeneity vs silhouette score for various numbers of prototypes.
  • E.5: study homogeneity vs k-means score for various numbers of prototypes.
  • E.6: computation time study.
  • E.7: Proofs of concepts with hyperbolic DR kernels.
E.1 Experimental setting
@@
Implementation. Throughout, the spectral clustering implementation of scikit-learn [56] is used to perform either the clustering steps or the initialization of transport plans. For all methods, ${\mathbf{Z}}$ is initialized from i.i.d. sampling of the standard Gaussian distribution $\mathcal{N}(0,1)$ and further optimized using PyTorch’s automatic differentiation [55] with Adam optimizer [35]. OT-based solvers are built upon the POT [27] library. k-Means is performed using the scikit-learn [56] implementation.
Validated hyperparameters. For the SEA based similarities, we validated perplexity across the set $\{20,50,100,150,200,250\}$. For all kernels, the number of output samples $n$ spans a set of $10$ values, starting at the number of classes in the data and incrementing in steps of $20$. For the computation of ${\mathbf{T}}$ in DistR (see Section 4.1), we benchmark our Conditional Gradient solver, and the Mirror Descent algorithm whose hyperparameter $\varepsilon$ is validated in the two first values within the set $\{10^{i}\}_{i=-3}^{3}$ leading to stable optimization.
Datasets. We provide details about the datasets used in our study. For image datasets, we use COIL-20333https://www1.cs.columbia.edu/CAVE/software/softlib/coil-100.php [53], MNIST and fashion-MNIST444taken from Torchvision [46]. [87]. Regarding single-cell genomics datasets, we rely on PBMC 3k555downloaded from Scanpy [86]. [86], SNAREseq [^6] https://github.com/rsinghlab/SCOT chromatin and gene expression [15] and the scRNA-seq dataset [^7] https://github.com/solevillar/scGeneFit-python from [90] with two hierarchical levels of label. Dimensions are provided in Table 1 When the dimensionality of a dataset exceeds $50$, we pre-process it by applying a PCA in dimension $50$, as done in practice [76].
@@
Scores. The homogeneity and NMI scores are taken from Torchmetrics [21]. The other scores used in the experiments are computed as follows.
i) Silhouette: The first step is to associate a label with each prototype ${\mathbf{z}}_{k}$ defined by the $y$ maximizing $\sum_{i\in{[\![N]\!]}}T_{ik}\bm{1}_{y_{i}=y}$ where $y_{i}$ is the class label of the input data point ${\mathbf{x}}_{i}$. Then, to incorporate the relative importance $w_{i}=\left[{\mathbf{h}}_{Z}\right]_{i}$ of each prototype ${\mathbf{z}}_{i}$, we define the weighted mean intra-cluster and nearest-cluster distances that read
$a_{i}({\mathbf{Z}},{\mathbf{Y}},{\mathbf{w}})=\frac{\sum_{j\in{[\![N]\!]}\delta_{y_{i}=y_{j}}w_{j}d({\mathbf{z}}_{j},{\mathbf{z}}_{j})}}{\sum_{j\in{[\![N]\!]}\delta_{y_{i}=y_{j}}w_{j}}}\quad\text{and}\quad b_{i}({\mathbf{Z}},{\mathbf{Y}},{\mathbf{w}})=\frac{\sum_{j\in{[\![N]\!]}}\delta_{y_{j}=k}w_{j}d({\mathbf{z}}_{j},{\mathbf{z}}_{i})}{\sum_{j\in{[\![N]\!]}}\delta_{y_{j}=k}w_{j}}$ (81)
such that the silhouette coefficient is $s_{i}({\mathbf{Z}},{\mathbf{Y}},{\mathbf{w}})=\frac{b_{i}-a_{i}}{\max(b_{i},a_{i})}$ and the final silhouette score reads as
$S({\mathbf{Z}},{\mathbf{Y}},{\mathbf{w}})=\sum_{i\in{[\![N]\!]}}w_{i}s_{i}({\mathbf{Z}},{\mathbf{Y}},{\mathbf{w}})\>.$ (82)
ii) k-means: We first run the k-means algorithm on the prototypes $\{{\mathbf{z}}_{k}\}_{k\in{[\![n]\!]}}$ giving us the predicted label $\tilde{y}_{k}$ for each $k\in{[\![n]\!]}$. The predicted label for each input data point $i\in{[\![N]\!]}$ is then given by its prototype’s predicted label i.e. $\hat{y}_{i}=\tilde{y}_{\arg\max_{k}T_{ik}}$. Then we compute the NMI score [38] between $(\hat{y}_{i})_{i\in{[\![N]\!]}}$ and the ground-truth class labels $(y_{i})_{i\in{[\![N]\!]}}$.
E.2 Comparison with COOT clustering
The CO-Optimal-Transport (COOT) problem, proposed in [60], is as follows,
@@ E.2 Comparison with COOT clustering
where ${\mathbf{h}}_{r}\in\Sigma_{N}$, $\overline{{\mathbf{h}}}_{r}\in\Sigma_{n}$, ${\mathbf{h}}_{c}\in\Sigma_{p}$ and $\overline{{\mathbf{h}}}_{c}\in\Sigma_{d}$. One can seek to optimize the above objective with respect to ${\mathbf{Z}}$ to obtain a competitor method to DistR. This problem is called COOT clustering in [60]. In the latter, ${\mathbf{T}}_{r}$ then plays the role of a soft clustering matrix of the rows of ${\mathbf{X}}$ while ${\mathbf{T}}_{c}$ can be seen as a soft clustering matrix of its columns. The above is thus a linear DR model.
@@
In Table 2, we display the homogeneity values obtained with COOT along with the methods described Section 5. Precisely, it measures to what extent the clustering given by ${\mathbf{T}}_{r}$ groups points with the same ground truth label. One can notice that COOT falls short compared to its competitors that leverage affinity matrices as in state of the art (non-linear) DR methods.
E.3 Complete scores
We complete the results shown in Section 5 by providing the scores obtained on all datasets and models. Scores are plotted in Figure 6 for the t-SNE model and in Figure 7 for the PCA model.
Figure 6: Scores (×$\times 100$) with respect to the number of prototypes (in $\mathbb{R}^{2}$) produced by DistR using the t-SNE model: SEA similarity for ${\mathbf{C}}_{X}$ [75], Student’s kernel for ${\mathbf{C}}_{Z}$ and loss $L_{\operatorname{KL}}$.
Figure 7: Scores (×$\times 100$) with respect to the number of prototypes (in $\mathbb{R}^{10}$) produced by DistR using the PCA model: $\langle,\rangle_{\mathbb{R}^{p}}$ similarity for ${\mathbf{C}}_{X}$ [75], $\langle,\rangle_{\mathbb{R}^{d}}$ for ${\mathbf{C}}_{Z}$ and loss $L_{2}$.
E.4 Dynamics between homogeneity and silhouette scores across numbers of prototypes
@@
E.5 Dynamics between homogeneity and NMI scores across numbers of prototypes
@@
E.6 Computation time comparison
We compare in the following the computation time for all methods benchmarked in Table 1 when using a spectral method and a SNE method. All experiments were done on a server using a GPU (Tesla V100-SXM2-32GB) and composed of 18 cores Intel(R) Xeon(R) Gold 6240 CPU @ 2.60GHz. All DR steps benefit from our GPU compatible implementation, while spectral clustering (SC) was performed on CPU using scikit-learn implementation running on CPUs.
@@ E.6 Computation time comparison
E.7 Proofs of concepts with hyperbolic kernels
Hyperbolic spaces [13, 26, 30, 41] are of particular interest as they can capture hierarchical structures more effectively than Euclidean spaces and mitigate the curse of dimensionality by producing representations with lower distortion rates. For instance, [30] adapted t-SNE by using the Poincaré distance and by changing the Student’s t-distribution with a more general hyperbolic Cauchy distribution. Notions of projection subspaces can also be adapted, e.g. [13] use horospheres as one-dimensional subspaces. To match our experiments with the neighbor embeddings in Euclidean settings, we adapt the Symmetric Entropic Affinity (SEA) from [75] for ${\mathbf{C}}_{X}$ and the scalar-normalized student similarity for ${\mathbf{C}}_{Z}$ [76], by simply changing the Euclidean distance by an hyperbolic distance.
@@ E.7 Proofs of concepts with hyperbolic kernels
Computations in Hyperbolic spaces are done with Geoopt [36] and the RAdam optimizer [3] replaces Adam. A Wrapped Normal distribution in Hyperbolic spaces [52] is used to initialize ${\mathbf{Z}}$ in the hyperbolic setting. All the computations were conducted in the Lorentz model [54], which is less prone to numerical errors. As such, we used the associated distance function to form ${\mathbf{C}}_{Z}$. After optimization, results are projected back to the Poincaré ball for visualization purpose. In this hyperbolic context, we adopted the formulation of [30] which generalize Student’s t-distributios by Hyperbolic Cauchy distributions (denoted as H-Student in the results). Notice that [30] considered weighted sums of DR objective depending respectively on $L_{2}$ and $L_{KL}$, including 2 additional hyperparamaters, plus various validated curvature levels for the inner hyperbolic distances. In the following experiments we only kept $L_{KL}$ for comparison with the Euclidean SNE based methods illustrated in Section 5 and previous Sections of E, while validating the same hyperparameters and setting the space curvature to $1$. The silhouette score was adapted to this kernel considering the Hyperbolic distance insteand of the euclidean one, and we implemented a Hyperbolic Kmeans whose barycenters are estimated using the RAdam optimizer to compute the NMI scores.

### v3

@@ Appendix F Appendix of experimental section
  • F.1: implementation details, validation of hyperparameters, datasets and metrics.
  • F.2: comparison with COOT clustering.
  • F.3: Best trade-off between metrics using t-SNE with $l_{2}$ symmetrization and UMAP.
  • F.4: complete scores on all datasets for all kernels relating to PCA, t-SNE and UMAP.
  • F.5: study homogeneity vs silhouette score for various numbers of prototypes.
  • F.6: study homogeneity vs k-means score for various numbers of prototypes.
  • F.7: Benchmark between Spectral and Kmeans clustering
  • F.8: study sensitivity w.r.t the embedding dimension $d$ on spectral methods.
  • F.9: computation time study.
  • F.10: Proofs of concepts with hyperbolic DR kernels.
F.1 Experimental setting
@@
Implementation. Throughout, the spectral clustering implementation of scikit-learn (Pedregosa et al., 2011) is used to perform either the clustering steps or the initialization of transport plans. For all methods, ${\mathbf{Z}}$ is initialized from i.i.d. sampling of the standard Gaussian distribution $\mathcal{N}(0,1)$ and further optimized using PyTorch’s automatic differentiation (Paszke et al., 2017) with Adam optimizer (Kingma and Ba, 2014). OT-based solvers are built upon the POT (Flamary et al., 2021) library. k-means is performed using the scikit-learn (Pedregosa et al., 2011) implementation.
Validated hyperparameters. For the SEA and UMAP based similarities, we validated perplexity across the set $\{20,50,100,150,200,250\}$. For UMAP, we further optimized before learning prototypes, the coefficients $a$ and $b$ involved in the parameterized Student kernel for ${\mathbf{C}}_{Z}$ using the algorithm provided by McInnes et al. (2018). For all kernels, the number of output samples $n$ spans a set of $10$ values, starting at the number of classes in the data and incrementing in steps of $20$. For the computation of ${\mathbf{T}}$ in DistR (see Section 4.1), we benchmark our Conditional Gradient solver, and the Mirror Descent algorithm whose hyperparameter $\varepsilon$ is validated in the two first values within the set $\{10^{i}\}_{i=-3}^{3}$ leading to stable optimization.
Datasets. We provide details about the datasets used in our study. For image datasets, we use COIL-20222https://www1.cs.columbia.edu/CAVE/software/softlib/coil-100.php (Nene et al., 1996), MNIST and fashion-MNIST333taken from Torchvision (Marcel and Rodriguez, 2010). (Xiao et al., 2017). Regarding single-cell genomics datasets, we rely on PBMC 3k444downloaded from Scanpy (Wolf et al., 2018). (Wolf et al., 2018), SNAREseq [^5] https://github.com/rsinghlab/SCOT chromatin and gene expression (Chen et al., 2019) and the scRNA-seq dataset [^6] https://github.com/solevillar/scGeneFit-python from (Zeisel et al., 2015) with two hierarchical levels of label. Dimensions are provided in Table 1 When the dimensionality of a dataset exceeds $50$, we pre-process it by applying a PCA in dimension $50$, as done in practice (Van der Maaten and Hinton, 2008).
@@
Scores. The homogeneity and NMI scores are taken from Torchmetrics (Detlefsen et al., 2022). The other scores used in the experiments are computed as follows.
i) Silhouette: The first step is to do a weighted majority vote and associate a label $\tilde{y}_{k}$ to each prototype ${\mathbf{z}}_{k}$. This label is defined by the $y$ maximizing $\sum_{i\in{[\![N]\!]}}T_{ik}\bm{1}_{y_{i}=y}$ where $y_{i}$ is the true class label of the input data point ${\mathbf{x}}_{i}$. Then, to incorporate the relative importance $w_{j}=\left[{\mathbf{h}}_{Z}\right]_{j}$ of each prototype ${\mathbf{z}}_{j}$, we define the weighted mean intra-cluster and nearest-cluster distances that read
$\forall k\in{[\![n]\!]},\ a_{k}({\mathbf{Z}},{\mathbf{Y}},{\mathbf{w}})=\frac{\sum_{j\in{[\![n]\!]}}\bm{1}_{\tilde{y}_{j}=\tilde{y}_{k}}w_{j}d({\mathbf{z}}_{j},{\mathbf{z}}_{k})}{\sum_{j\in{[\![n]\!]}}\bm{1}_{\tilde{y}_{j}=\tilde{y}_{k}}w_{j}}\quad\text{and}\quad b_{k}({\mathbf{Z}},{\mathbf{Y}},{\mathbf{w}})=\min_{k^{\prime}\neq k}\frac{\sum_{j\in{[\![n]\!]}}\bm{1}_{\tilde{y}_{j}=k^{\prime}}w_{j}d({\mathbf{z}}_{j},{\mathbf{z}}_{k})}{\sum_{j\in{[\![n]\!]}}\bm{1}_{\tilde{y}_{j}=k^{\prime}}w_{j}}$ (81)
such that the silhouette coefficient is $s_{k}({\mathbf{Z}},{\mathbf{Y}},{\mathbf{w}})=\frac{b_{k}-a_{k}}{\max(b_{k},a_{k})}$ and the final silhouette score reads as
$S({\mathbf{Z}},{\mathbf{Y}},{\mathbf{w}})=\sum_{k\in{[\![n]\!]}}w_{k}s_{k}({\mathbf{Z}},{\mathbf{Y}},{\mathbf{w}})\>.$ (82)
ii) k-means: We first run the k-means algorithm on the prototypes $\{{\mathbf{z}}_{k}\}_{k\in{[\![n]\!]}}$ giving us the predicted label $\tilde{y}_{k}$ for each $k\in{[\![n]\!]}$. The predicted label for each input data point $i\in{[\![N]\!]}$ is then given by its prototype’s predicted label i.e. $\hat{y}_{i}=\tilde{y}_{\arg\max_{k}T_{ik}}$. Then we compute the NMI score (Kvålseth, 2017) between $(\hat{y}_{i})_{i\in{[\![N]\!]}}$ and the ground-truth class labels $(y_{i})_{i\in{[\![N]\!]}}$.
F.2 Comparison with COOT clustering
The CO-Optimal-Transport (COOT) problem, proposed in (Redko et al., 2020), is as follows,
@@ F.2 Comparison with COOT clustering
where ${\mathbf{h}}_{r}\in\Sigma_{N}$, $\overline{{\mathbf{h}}}_{r}\in\Sigma_{n}$, ${\mathbf{h}}_{c}\in\Sigma_{p}$ and $\overline{{\mathbf{h}}}_{c}\in\Sigma_{d}$. One can seek to optimize the above objective with respect to ${\mathbf{Z}}$ to obtain a competitor method to DistR. This problem is called COOT clustering in Redko et al. (2020). In the latter, ${\mathbf{T}}_{r}$ then plays the role of a soft clustering matrix of the rows of ${\mathbf{X}}$ while ${\mathbf{T}}_{c}$ can be seen as a soft clustering matrix of its columns. The above is thus a linear DR model.
@@
In Table 2, we display the homogeneity values obtained with COOT along with the methods described Section 5. Precisely, it measures to what extent the clustering given by ${\mathbf{T}}_{r}$ groups points with the same ground truth label. One can notice that COOT falls short compared to its competitors that leverage affinity matrices as in state-of-the-art (non-linear) DR methods.
F.3 Best trade-off between metrics using t-SNE and UMAP
Figure 6: Complements to Figure 3 of the main paper for t-SNE with $l_{2}$ symmetrization and UMAP : Best trade-off between homogeneity vs silhouette (2 first plots), and homogeneity vs NMI (2 last plots). Scores are normalized in $\left[0,1\right]$ via min-max scaling over a dataset. Small markers are scores per dataset for 5 runs, while big ones are their mean. We illustrate the 20-80% percentiles of scores per method as a colored surface.
F.4 Complete scores
We complete the results shown in Section 5 by providing the scores obtained on all datasets and models. Scores are plotted in Figure 7 for the PCA model, in Figure 8 for the t-SNE model with entropic symmetrization, then in Figure 9 with $l_{2}$ symmetrization, and Figure 10 for the UMAP model.
Figure 7: Scores (×$\times 100$) with respect to the number of prototypes (in $\mathbb{R}^{10}$) produced by DistR using the PCA model: $\langle,\rangle_{\mathbb{R}^{p}}$ similarity for ${\mathbf{C}}_{X}$ (Van Assel et al., 2023), $\langle,\rangle_{\mathbb{R}^{d}}$ for ${\mathbf{C}}_{Z}$ and loss $L_{2}$.
Figure 8: Scores (×$\times 100$) with respect to the number of prototypes (in $\mathbb{R}^{2}$) produced by DistR using the t-SNE model: SEA similarity for ${\mathbf{C}}_{X}$ (Van Assel et al., 2023), Student’s kernel for ${\mathbf{C}}_{Z}$ and loss $L_{\operatorname{KL}}$.
Figure 9: Scores (×$\times 100$) with respect to the number of prototypes (in $\mathbb{R}^{2}$) produced by DistR using the t-SNE model with $l_{2}$ symmetrization: EA similarity for ${\mathbf{C}}_{X}$ (Van der Maaten and Hinton, 2008), Student’s kernel for ${\mathbf{C}}_{Z}$ and loss $L_{\operatorname{KL}}$.
Figure 10: Scores (×$\times 100$) with respect to the number of prototypes (in $\mathbb{R}^{2}$) produced by DistR using the UMAP model: t-conorm of the smoothed nearest neighbors $v_{j|i}=e^{-\|{\mathbf{x}}_{i}-{\mathbf{x}}_{j}\|_{2}-\rho_{i}/\sigma_{i}}$ for ${\mathbf{C}}_{X}$ (McInnes et al., 2018, Eq.16), parameterized Student’s kernel for ${\mathbf{C}}_{Z}$ (McInnes et al., 2018, Eq.17) and loss $L_{BCE}$.
F.5 Dynamics between homogeneity and silhouette scores across numbers of prototypes
@@
F.6 Dynamics between homogeneity and NMI scores across numbers of prototypes
@@
F.7 Benchmark between Spectral and Kmeans clustering
[…]
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 15: Best trade-off between homogeneity vs silhouette (left), and homogeneity vs NMI (right). Scores are normalized in $\left[0,1\right]$ via min-max scaling over a dataset. Small markers represent scores averaged over 5 runs for a given dataset, while big ones are their mean over datasets. The illustration follows the same principal than fig. 3
F.8 Sensitivity analysis w.r.t the embedding dimension
Figure 16: Sensitivity analysis w.r.t the embedding dimension $d$ for the spectral method analog to PCA with $C_{X}({\mathbf{X}})={\mathbf{X}}{\mathbf{X}}^{\top}$ and $C_{Z}({\mathbf{Z}})={\mathbf{Z}}{\mathbf{Z}}^{\top}$. Over three datasets, we make vary the embedding dimension from $d\in\{12,14,...,20\}$.
F.9 Computation time comparison
We compare in the following the computation time for all methods benchmarked in Table 1 when using a spectral method and a SNE method. For fair comparison across methods, we do not apply here the low-rank factorizations used for experiments of the main paper. All experiments were done on a server using a GPU (Tesla V100-SXM2-32GB) and composed of 18 cores Intel(R) Xeon(R) Gold 6240 CPU @ 2.60GHz. All DR steps benefit from our GPU compatible implementation, while spectral clustering (SC) was performed on CPU using scikit-learn implementation running on CPUs.
@@ F.9 Computation time comparison
F.10 Proofs of concepts with hyperbolic kernels
Hyperbolic spaces (Chami et al., 2021, Fan et al., 2022, Guo et al., 2022, Lin et al., 2023) are of particular interest as they can capture hierarchical structures more effectively than Euclidean spaces and mitigate the curse of dimensionality by producing representations with lower distortion rates. For instance, Guo et al. (2022) adapted t-SNE by using the Poincaré distance and by changing the Student’s t-distribution with a more general hyperbolic Cauchy distribution. Notions of projection subspaces can also be adapted, e.g. Chami et al. (2021) use horospheres as one-dimensional subspaces. To match our experiments with the neighbor embeddings in Euclidean settings, we adapt the Symmetric Entropic Affinity (SEA) from Van Assel et al. (2023) for ${\mathbf{C}}_{X}$ and the scalar-normalized student similarity for ${\mathbf{C}}_{Z}$ (Van der Maaten and Hinton, 2008), by simply changing the Euclidean distance by an hyperbolic distance.
@@ F.10 Proofs of concepts with hyperbolic kernels
Computations in Hyperbolic spaces are done with Geoopt (Kochurov et al., 2020) and the RAdam optimizer (Bécigneul and Ganea, 2018) replaces Adam. A Wrapped Normal distribution in Hyperbolic spaces (Nagano et al., 2019) is used to initialize ${\mathbf{Z}}$ in the hyperbolic setting. All the computations were conducted in the Lorentz model (Nickel and Kiela, 2018), which is less prone to numerical errors. We used the distance function from Nickel and Kiela (2018) to form ${\mathbf{C}}_{Z}$. Specifically, we employed a scalar-normalized Gaussian kernel, where the distances between points are computed as in Nickel and Kiela (2018, Equation 1). After optimization, results are projected back to the Poincaré ball for visualization purposes. In this hyperbolic context, we adopted the formulation of Guo et al. (2022) which generalizes Student’s t-distribution by Hyperbolic Cauchy distributions (denoted as H-Student in the results). Notice that Guo et al. (2022) considered weighted sums of DR objective depending respectively on $L_{2}$ and $L_{KL}$, including 2 additional hyperparamaters, plus various validated curvature levels for the inner hyperbolic distances. In the following experiments, we only kept $L_{KL}$ for comparison with the Euclidean SNE-based methods illustrated in Section 5 and previous Sections of F, while validating the same hyperparameters and setting the space curvature to $1$. The silhouette score was adapted to this kernel considering the Hyperbolic distance instead of the Euclidean one, and we implemented a Hyperbolic Kmeans whose barycenters are estimated using the RAdam optimizer to compute the NMI scores.